OpenStudy (anonymous):

2+(x-1)(x+1)=x(x-2) Solve for x

OpenStudy (anonymous):

2+(x-1)(x+1)=x^2-2x 2+(x^2-1)=x^2-2x 2+x^2-1=x^2-2x x^2+2-1=x^2-2x x^2+1=x^2-2x x^2-x^2-2x+1=0 -2x+1=0 x=-1/(-2)=1/2

OpenStudy (anonymous):

first just factor it all out \[2+x ^{2}+2x+1=x ^{2}-2x\] \[2+1=-4x\] x=-(3/4)

OpenStudy (anonymous):

paul (x-1)(x+1)=x^2-1

OpenStudy (anonymous):

sorry my mistake he's right!

OpenStudy (anonymous):

okay, so its 1/2?